OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2)

OpenStudy (anonymous):

can you write the equation in y=mx+c form?

OpenStudy (anonymous):

I don't remember how to do this at all.

OpenStudy (anonymous):

isolate y on left hand side and all the other things on right!

OpenStudy (anonymous):

can you make up the equation now?

OpenStudy (anonymous):

No :/

OpenStudy (anonymous):

ok.3x-y=-3 -y=-3-3x y=3+3x will be the equation.can you determine the slope from the equation?

OpenStudy (anonymous):

slope is the number with x which in this case 3,right?

OpenStudy (anonymous):

y = 3x − 14 < Answer (: Thanksssss!

OpenStudy (anonymous):

great.you did it!